OpenStudy (anonymous):

Cathy did a survey in four bird parks and wrote the following observations: Bird Park Survey Park Number of Green Birds Total Number of Birds P 18 82 Q 39 88 R 38 64 S 22 50 Based on the survey, which park had the greatest percentage of green birds?

OpenStudy (anonymous):

@HazelLuv99

OpenStudy (anonymous):

help?

OpenStudy (shaik0124):

just a min i will help u

OpenStudy (hazelluv99):

ME NO GOOD AT THIS I SOWWY

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

lol hazel grace

OpenStudy (igreen):

Find the percentages.

OpenStudy (shaik0124):

for calculating percentage number of green birds/total number of birds then multiply with 100

OpenStudy (igreen):

18 / 82 * 100 = ? 39 / 88 * 100 = ? 38 / 64 * 100 = ? 22 / 50 * 100 = ?

OpenStudy (shaik0124):

for case 1 it is (18/82) *100

OpenStudy (anonymous):

P 22% Q 44% R 59% S 44%

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

So Park R had the most green birds.

OpenStudy (anonymous):

ok good

OpenStudy (shaik0124):

59% is answer

OpenStudy (anonymous):

thank you

OpenStudy (igreen):

No problem, Happy to Help \(\Huge\ddot\smile\)